OpenStudy (anonymous):

Suppose that you have $114 for tickets, and you want 6 tickets in the upper box section for weekday and weekend games. A ticket for the weekday is $18 and the weekend is $21. How many of each ticket can you buy?

OpenStudy (anonymous):

I got the answer, x=4 y=2, but I don't know how to show my work

OpenStudy (mathstudent55):

Let x = number of $18 ticket Let y = number of $21 ticket x + y = 6 the total number of tickets is 6 18x + 21y = 114 18x is the cost of x $18-tickets; 21y is the cost of y $21-tickets; 18x + 21y is the total spent = 114 x + y = 6 18x + 21y = 114 Solve the system of equations.

OpenStudy (anonymous):

y=2 x=4

OpenStudy (mathstudent55):

You already had written you knew the answer. Do you know how to solve the system of equations, or do you need help with it?

OpenStudy (anonymous):

I need help with solving the system of equations (showing my work)

OpenStudy (anonymous):

I've gotten to the step 18(6-y)+21y=114... after that I am stuck

OpenStudy (mathstudent55):

Ok x + y = 6 18x + 21y = 114 Let's solve the system of equations using the substitution method. In this method, you first solve one equation for one variable. Oh, I see, you're getting there. x = 6 - y 18(6 - y) + 21y = 114 This is exactly what you have so far.

OpenStudy (mathstudent55):

Now distribute the 18

OpenStudy (anonymous):

108-18y+21y=114

OpenStudy (mathstudent55):

18(6) - 18(y) + 21y = 114

OpenStudy (mathstudent55):

Right

OpenStudy (mathstudent55):

Now combine like terms on the left side

OpenStudy (anonymous):

108-39y=114 ?

OpenStudy (mathstudent55):

Careful with that addition. -18y + 21y You are adding two quantities with different signs, -18y and 21y

OpenStudy (anonymous):

ohhhhhhh okkk! Thank you for catching that!!! 108+3y=114

OpenStudy (anonymous):

and then -108 from both sides ???

OpenStudy (anonymous):

3y=6/3 y=2!

OpenStudy (mathstudent55):

The rule is take absolute values of the numbers: |21| = 21 |-18| = 18 Then subtract the smaller one from the larger one: 21 - 18 = 3 Then the sign is the same sign as the number that has the larger absolute value. 21 has a larger abs value than -18, and 21 is positive, so the answer is positive. so -18 + 21 = +3

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

3y = 6 y = 2. Correct. Now substitute y = 2 into one of the original equations to get x.

OpenStudy (anonymous):

x=4

OpenStudy (mathstudent55):

x + 2 = 6 x = 4 Correct. Answer is: 4 $18 tickets and 2 $21 tickets. Notice there is no x and y in the answer above. That is because the question does not use any x and y. We used x and y to represent the question mathematically. The person asking the question wants to know the numbers of tickets and has no idea what x and y are.

OpenStudy (anonymous):

ahh right. Your attention to detail saved me! haha thanks so much!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

What grade are you in? (math student)